John kicks the ball, and the ball does a projectile motion with an angle of 53o to horizontal with an initial velocity of 10 m/s. What is (a) its maximum height, (b) its maximum distance, and (c) its hangtime?


1
Expert's answer
2021-10-10T15:56:21-0400

(a)

"H_{max} = \frac{u^2sin^2 θ}{2g} \\

= \frac{10^2 \times sin^2 53}{2 \times 9.81} \\

= 3.25086 \;m"

(b)

"R_{max} = \frac{u^2 sin2θ }{g} \\

= \frac{10^2 sin(2 \times 53)}{9.81} \\

= 9.7988 \;m"

(c)

"T= \frac{2u sin θ}{g} \\

= \frac{2 \times 10 \times sin53}{9.81} \\

= 1.6282 \;s"